OpenStudy (anonymous):

how would i go about solving 8e^2x=20. btw we are on logs.

OpenStudy (anonymous):

e^2x=20/8 lne^2x = ln (20/8) 2x = ln (20/8) x = [ln(20/80]/2

OpenStudy (anonymous):

\[8e^{2x}=20\] \[e^{2x}=4\]\[2x=ln(4)\] \[x=\frac{ln(4)}{2}\]

OpenStudy (anonymous):

20/8 = 4 :-??

OpenStudy (anonymous):

and since \[ln(4) = ln(2^2)=2ln(2)\] \[\frac{ln(4)}{2}=ln(2)\]

OpenStudy (anonymous):

ho ho ho

OpenStudy (anonymous):

time for me to retire

OpenStudy (anonymous):

;-)

OpenStudy (anonymous):

\[\frac{ln(5)}{2}\]

OpenStudy (anonymous):

or whatever \[\frac{20}{8}\] is

OpenStudy (anonymous):

they already answer it in the other question box :D

OpenStudy (anonymous):

hope they didn't write \[\frac{ln(4)}{2}\]!

OpenStudy (anonymous):

haha